### v1

# Implicit Neural Representations for Spatial Transcriptomics

Abstract
Spatial Transcriptomics (ST) is a method that captures spatial gene expression profiles within histological sections. The discrete spatial distribution and the super-high dimensional sequencing results make ST data challenging to be modeled effectively. In this paper, we manage to model ST in a continuous and compact manner by the proposed tool, SUICA, empowered by the great approximation capability of Implicit Neural Representations (INRs) that can improve both the spatial resolution and the gene expression. Concretely within the proposed SUICA, we incorporate a graph-augmented Autoencoder to effectively model the context information of the unstructured spots and provide informative embeddings that are structure-aware for spatial mapping. We also tackle the extremely skewed distribution in a regression-by-classification fashion and enforce classification-based loss functions for the optimization of SUICA. By extensive experiments of a wide range of common ST platforms, SUICA outperforms both conventional INR variants and SOTA methods for ST super-resolution regarding numerical fidelity, statistical correlation, and bio-conservation. The prediction by SUICA also showcases amplified gene signatures that enriches the bio-conservation of the raw data and benefits subsequent analysis. The code is available at https://github.com/Szym29/SUICA.
1 Introduction
Spatial Transcriptomics (ST) enables scientists to quantify gene expression while preserving spatial information in tissue sections [27]. These platforms use various strategies to capture mRNA transcripts from tissue sections and perform sequencing to quantify the gene expressions at spatially defined locations [31]. However, achieving higher resolution is essential for accurately modeling and analyzing cellular functions [54], while for now, ST data can still be rather expensive (\$3,500/$cm^{2}$ for the capture chip and \$800/$cm^{2}$ for the high-density sequencing [2]). Beyond the cost constraints to get high-resolution ST, another key technical challenge lies in the drop-out rates (high sparsity of ST data) that makes the gene expressions less informative for analysis [42]. Compounding this challenge, different sequencing techniques and platforms exhibit significant heterogeneity in terms of spatial distribution, sequencing depth (number of mRNA readouts per spot), and drop-out rates [31]. As a result, such the difference makes a general backbone to model and analyze data across different tissues by different platforms less feasible. To address these technical challenges, we propose a computational framework that transforms discrete ST data into a continuous, compact representation. Our approach enables super-resolution spatial transcriptomics analysis, providing a robust foundation for downstream analyses.
Figure 1: Starting with the discretely sampled spots (a) of ST, SUICA performs continuous modeling (b) by aid of the great approximation power of INRs. This approach enables more precise profiling of cell heterogeneity, as visualized in the UMAP (c), which further facilitating the discovery of new biology.
Recently, Implicit Neural Representations (INRs) have drawn great attention from researchers for their compact, continuous, and differentiable properties as a novel representation of general coordinate-based signals. INRs map coordinates to corresponding values with a neural network with wide applications in inverse graphics [29], geometric modeling [38], and video compression [4]. Motivated by INRs’ inherent smoothness property, we leverage this characteristic to interpolate between sparse ST spots, enabling more comprehensive spatial transcriptomics analysis. However, applying INRs to ST data faces two major challenges. First, scaling INRs to the super-high dimensional space of gene expression is non-trivial, as simply widening or deepening the network cannot effectively overcome the curse of dimensionality. Second, the low mRNA transcript capture rate in ST, combined with varying gene expression patterns across cellular states, results in zero-inflated ST data with high sparsity [41]. This sparsity makes it particularly challenging for conventional INRs to accurately capture the underlying complex, non-linear spatial patterns.
Building upon this approach, we introduce SUICA, a powerful variant of INR designed especially for ST data. SUICA fully accounts for the unique properties of ST and the diversity of current mainstream ST platforms. In SUICA, we employ an Autoencoder (AE) based on Graph Convolutional Network (GCN) [18] to bridge the gap between conventional INRs and ST. The GCN enhances the approximation capability for unstructured data by leveraging contextual information and making the embeddings structure-aware. By leveraging the strong profiling power of Graph Autoencoder (GAE) regarding the sparse and skewed distribution, we perform the fitting of INRs at the expressive low-dimensional embedding space, which has been proven to be compact and more suitable for INRs. This approach facilitates more accurate and efficient modeling of ST data.
To optimize SUICA and enforce the sparsity within the regression-based scheme of INRs, we construct pseudo-probabilities and adopt a regression-by-classification approach for training. With SUICA, we are enabled to model diverse ST data from several popular platforms and sequencing protocols using a unified representation, facilitating comprehensive analysis as illustrated in Fig. 1. Extensive experimental results demonstrate that SUICA outperforms existing INR variants as well as the state-of-the-art methods in both numerical fidelity and statistical correlation. Moreover, SUICA’s imputation ability leads to more cell-type informed clustering results of ST. To summarize, our contributions in this paper are three-fold as follows:
  • We introduce SUICA, modeling ST data as a continuous and compact representation while preserving data authenticity;
  • We address the issue that prevent INRs from scaling to the super-high dimensional gene expression of ST by leveraging Graph Autoencoder and a classification-based loss function;
  • Extensive experiments show that SUICA achieves superior reconstruction quality and imputation capabilities on various ST datasets, facilitating subsequent analyses.
2 Related Work
2.1 Implicit Neural Representations
INRs model signals by mapping input coordinates to corresponding signal values using neural networks. Unlike conventional discrete grid-based signal representations, e.g., images, videos, and voxels, INRs are known for continuous modeling, allowing queries at arbitrary locations within the definition domain. Typically implemented as Multi-Layer Perceptrons (MLPs), INRs leverage the smoothness bias of MLPs to provide a certain level of interpolability, allowing effective generalization to unseen coordinates.
The idea of neural networks for function approximation dates back to the 1980s and 1990s [6, 14] as universal function approximation theories. From a relatively modern perspective, SIREN [43] adopts periodic sine functions as the activation function and learns INRs rich in fine details. Equipped with Random Fourier Features, FFN [48] multiplies the coordinates with a random feature sampled from a normal distribution. BACON [23] further explores the spectral bandwidth of the target signal and progressively regresses in a coarse-to-fine manner.
INRs have been applied across various downstream tasks, including inverse graphics and video compression. NeRF [29] parameterizes the 3D scene as an MLP-based radiance field with colors and densities. It applies differential volume rendering to query a series of points in the field and then calculate the weighted sum as the pixel value. Following this parameterization, subsequent works [24, 56, 46, 16, 11, 32, 40, 1, 3] further improve the radiance field for efficient training and rendering. Additionally, the continuity of INR is well-suited for application in 3D geometric reconstruction methods [38, 52, 55, 33]. As for video compression, NeRV [4] and subsequent variants [22, 5] employ compact INRs to reduce data redundancy.
It is noteworthy that existing INR applications map from low-dimensional to low-dimensional spaces, e.g., $\mathbb{R}^{2}\to\mathbb{R}^{3}$ for image regression, and $\mathbb{R}^{5}\to\mathbb{R}^{4}$ for inverse graphics. The use of INRs for mapping from low dimension (i.e., $\mathbb{R}^{2}$) to super-high dimension (i.e., over 20,000 channels) remains unexplored.
Figure 2: The overall pipeline of SUICA. SUICA is a powerful tool powered by INRs to improve both spatial resolution and cell profiles (gene expression).
2.2 Deep Learning for Spatial Transcriptomics
Deep learning methods are increasingly adopted in ST to enhance spatial resolution, denoise raw data, improve spatial domain clustering, and support more accurate downstream analysis [57]. Given the complexity of capturing cell-to-cell and spatial dependencies, graph-based methods leveraging cell neighborhood relationships have gained significant traction. Notable examples include GCN-based methods like SpaGCN [15], STAGATE [9], and GraphST [25], as well as graph transformer-based model like SiGra [49]. These graph-based methods progressively refine the modeling of ST data, enabling more nuanced interpretation of complex tissue structure.
ST data often require a higher resolution to accurately capture fine spatial patterns within tissues. Several methods leverage histological images to enhance the spatial resolution of gene expression data, including Convolution Neural Network (CNN)-based models ST-Net [12] and DeepSpaCE [30], and Vision Transformer-based methods such as HistoGene [37] and Hist2ST [58]. In cases without histological images, other approaches  [59, 61, 21] rely solely on spatial and gene expression data to improve resolution. To generate unmeasured spots profiles, STAGE [21] trains a spatial location-supervised Autoencoder that enhances resolution by integrating spatial coordinates with gene expression data. In summary, these resolution enhancement methods significantly improve ST quality, improving spatial visualization and facilitating more accurate downstream analyses.
3 SUICA
3.1 Preliminary
3.1.1 Spatial Transcriptomics
ST data can be viewed as a point cloud, where each point is parameterized by spatial coordinates $\mathbf{x}$. The mRNA readouts are bounded at these sampled locations and represented by a high dimensional vector, where each channel reflects the numerical abundance of a particular mRNA transcript (0 indicates absence or failed capture).
In terms of the spatial distribution of the sampled spots, different ST platforms follow different sampling protocols, e.g., Visium applies a fixed array of probes so the sampled points are distributed uniformly over the whole slice, whereas other common platforms do not guarantee such regularity. Thus, to accommodate general use cases, ST data is often modeled as unstructured without quantizing into regular grids, making CNN-based analysis inapplicable.
3.1.2 Challenges
The key challenge in modeling ST lies in the super-high dimensional representation, often exceeding 20,000 channels. Due to low mRNA capture rates and varying gene expression patterns across cellular states, ST data is typically highly sparse, with zero values comprising up to 90%. A satisfying model must maintain the inherent sparsity of ST data, as it reflects the true gene expression. However, vanilla INRs are observed to have the tendency to yield normally distributed outputs that are smooth [20], rather than the zero-inflated ones of ST, presenting a significant challenge.
Beyond maintaining sparsity, it is also crucial to ensure the numerical fidelity of non-zero values for accurate cell type identification. It is worth noting that high sparsity complicates evaluation of reconstruction quality, as an entirely empty prediction may still result in low loss, yet be entirely unacceptable.
To this end, the design of the representations and evaluation protocol must address both sparsity and numerical fidelity—an aspect that has, to our knowledge, not been fully explored in the context of INRs.
3.2 Method
3.2.1 Overview
The overall pipeline is illustrated in Fig. 2. To construct a compact and dense embedding domain for subsequent INRs, we first incorporate a graph-based encoder and pre-train Graph Autoencoder (GAE) using the given ST slice in a self-regressing manner. With the pre-trained GAE, we obtain the encoded latent representation for all spots, denoted as $\mathbf{z}_{\text{gt}}$.
We then initiate INR tuning to optimize the neural mapping from ST coordinates to the embeddings, targeting $\mathbf{z}_{\text{gt}}$. Once the INR reaches a stable state, we attach a decoder that is pre-trained in our GAE. Subsequently, we fix the INR and train the decoder to fit the raw dataset readouts, $\mathbf{y}_{\text{gt}}$.
3.2.2 Graph-augmented Autoencoder
The sequenced readouts of ST are known for an extremely skewed distribution, represented as high-dimensional, sparse data, making them challenging for vanilla INRs to model effectively. This sparsity exacerbates the curse of dimensionality: as dimensionality increases, the volume of the data space expands rapidly, leading to sparsely distributed data. Consequently, the data required for reliable results grows exponentially with dimensionality, rendering data points increasingly dissimilar and challenging to organize efficiently.
To address this, we leverage an AE to transform the high-dimensional raw space into a compact, dense, and informative embedding space. Given the irregular structure, the context information, context or neighborhood information is integrated through a graph structure, which we model using a Graph Convolutional Network (GCN). Importantly, the GCN is incorporated solely in the encoder, as the decoder—responsible for reconstructing interpolated embeddings back into the raw space—lacks the requisite graph structure for such integration.
Figure 3: Spectral analysis with the embeddings of AE and GAE. GAE yields structure-aware and disentangled embeddings with high-frequency details. GTV: Graph Total Variation.
As illustrated in Fig. 3, we conduct a graph-based spectral analysis of embeddings generated by both the AE and the GAE. Specifically, we construct a connectivity graph using Euclidean distance between spots ($k=4$) and calculate the Graph Total Variation (GTV) for embeddings defined on the built graph. Fig. 3 is colored based on aggregated edge-wise variations of each vertex, revealing that the GAE produces shaper structural clues. Additionally, we compute the channel-wise variance for both embeddings. Together with the GTV results, these findings demonstrate that GAE offers a more expressive and informative representation, effectively capturing disentangled embeddings for ST.
To train GAE, we adopt the conventional Mean Square Error (MSE) for supervision as
$\mathcal{L}_{\text{gae}}=\frac{1}{|\mathbf{M}_{\mathbf{y}}|}\sum_{\mathbf{M}_{\mathbf{y}}}(\hat{\mathbf{y}}-\mathbf{y}_{\text{gt}})^{2},$ (1)
where $\mathbf{M}_{\mathbf{y}}$ represents the binary mask over $\mathbf{y}$.
3.2.3 Embedding Mapping
With the pre-trained GAE and encoded representation $\mathbf{z}_{\text{gt}}$ in place, we proceed to train the INR to map $\mathbf{x}\to\mathbf{z}$. The compact, dense embeddings provided by the GAE significantly reduce the workload of the subsequent INR, which must fit an otherwise highly skewed distribution.
For SUICA, we empirically select two baseline architectures: FFN [48] and SIREN [43], both of which have demonstrated effectiveness across various data types. In our experiments, we apply SIREN for spatially sparse ST and FFN for denser spatial distributions. A detailed discussion on this selection is included in the supplementary material.
Following conventional INR approaches, we assume Gaussian-distributed error in the predicted $\hat{\mathbf{z}}$ and use MSE loss for optimization:
$\mathcal{L}_{\text{embd}}=\frac{1}{|\mathbf{M}_{\mathbf{z}}|}\sum_{\mathbf{M}_{\mathbf{z}}}(\hat{\mathbf{z}}-\mathbf{z}_{\text{gt}})^{2},$ (2)
where $\mathbf{M}_{\mathbf{z}}$ represents the binary mask over elements in $\mathbf{z}$. While $\hat{\mathbf{z}}$ can serve as the mapping output for downstream analysis, we consider it as an intermediate result, decoding this latent code back into the raw space to achieve a more comprehensive reconstruction.
3.2.4 Decoding Head
A straightforward approach to decoding $\hat{\mathbf{z}}$ back to raw representations is to directly use the pre-trained GAE decoder for end-to-end INR tuning. However, this method encounters two primary challenges. (1) The embedding mapping of $\mathbf{x}\to\mathbf{z}$ is error-prone, and domain shifts can severely impair decoding performance, compounding with the inherent imperfections in GAE’s reconstruction. (2) A pre-trained decoder can become trapped in local minima, impeding the optimization of the embedding mapping if the INR depends on gradients from the decoder.
To address (1)(2), we first warm up the INR with Eq. 2 alone to stabilize the mapping, then attach the pre-trained decoder to learn the mapping from $\mathbf{z}\to\mathbf{y}$, keeping the INR fixed. This decoder-only training phase is designed to finetune a case-specific decoder that compensates for mapping errors and minimizes cumulative errors.
In INR regression tasks, norm-based loss functions (e.g., $\ell^{2}$ norm or MSE) are typically used, as they assume normally (Gaussian) distributed errors. However, for zero-inflated ST data, this assumption is invalid. To handle the imbalanced distribution of zero and non-zero values, we apply Dice Loss [45], which is sensitive to class imbalance and treats the regression task as a quasi-classification one. Dice Loss optimizes for Intersection over Union (IoU) between the prediction map and binary ground truth. Specifically, we use the non-negative half of $\tanh(\cdot)$ to map network outputs into a pseudo probability into a pseudo-probability range $[0,1)$, and compute the intersection using element-wise Hadamard products with the ground truth. To avoid division by 0, Dice Loss is computed as:
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum(\tanh(\hat{\mathbf{y}})\circ\text{sgn}(\mathbf{y}_{\text{gt}}))+\epsilon}{\sum\tanh(\hat{\mathbf{y}})+\sum\text{sgn}(\mathbf{y}_{\text{gt}})+\epsilon},$ (3)
where $\text{sgn}(\cdot)$ denotes the sign function returning 0 when the input is 0 and +1/-1 when the input is positive/negative. The full reconstruction loss function is then defined as:
$\mathcal{L}_{\text{recons}}=\frac{1}{|\mathbf{M}^{+}_{\mathbf{y}}|}\sum_{\mathbf{M}^{+}_{\mathbf{y}}}(\hat{\mathbf{y}}-\mathbf{y}_{\text{gt}})^{2}+\frac{1}{|\mathbf{M}_{\mathbf{y}}|}\sum_{\mathbf{M}_{\mathbf{y}}}|\hat{\mathbf{y}}-\mathbf{y}_{\text{gt}}|+\lambda\mathcal{L}_{\text{dice}},$ (4)
where $\mathbf{M}^{+}_{\mathbf{y}}$ represents the binary mask for $\mathbf{y}_{\text{gt}}>0$, and $\lambda$ mitigates Dice Loss’s numerical instability.
4 Experiments
4.1 Datasets & Metrics
For a quantitative benchmarking, we involve a nanoscale resolution Stereo-seq (SpaTial Enhanced REsolution Omics-sequencing) dataset, MOSTA [2]. MOSTA consists of a total of 53 sagittal sections from C57BL/6 mouse embryos at 8 progressive stages using Stereo-seq, from which we take 1 slice for each stage (from E9.5 to E16.5) for benchmarking. In addition to Stereo-seq, we also leverage ST data by two other common platforms, including Slide-seqV2 and 10x Genomics Visium, to demonstrate the generalization of SUICA.
As for evaluation of the fitting performance of SUICA with super-high dimensional data, we focus on 3 aspects, namely numerical fidelity, statistical correlation, and bio-conservation. For numerical fidelity, we apply MSE, MAE (Mean Absolute Error) and cosine similarity to measure the significant, subtle and directional errors between the predicted and ground-truth values respectively. Note that we only measure numerical fidelity on non-zero values considering the zero-inflated distribution of ST. To measure the statistical correlation, we employ Pearson Correlation Coefficient and Spearman’s Rank Correlation Coefficient (Spearman’s $\rho$), with both of them ranging from -1 to 1. Lastly, to evaluate how well the prediction preserves cellular heterogeneity and spatial coherence within the microenvironment of the slice, we use the Adjusted Rand Index (ARI) as a metric for quantifying bio-conservation leveraging the independent hand-crafted cell type annotations.
4.2 Implemented Details
We implement SUICA with PyTorch, sampling 80% of spots within a ST slice for training and retaining the remaining 20% for evaluation, under the evaluation protocol of super-resolution. Notably, this sampling scheme may place some validation spots outside the convex hull of the training set, making such interpolation challenging. To construct the $k$-NN graph for GCN, we set $k=5$, including the given cell itself. During the pre-training phase of GAE, we use Adam with a learning rate of $1e^{-5}$ for 200 epochs. After obtaining low-dimensional cell embeddings, we train the INR with Adam at a learning rate of $1e^{-4}$ for $1k$ epochs to learn the embedding mapping. Subsequently, the INR is frozen, with the pretrained decoder trained for an additional $1k$ epochs using Adam with the same learning rate. The deep learning-based reference-free cell super-resolution model, STAGE [21], is trained till convergence for benchmarking. All experiments are conducted on 1 NVIDIA RTX 4090.
4.3 Benchmarking Results
Stereo-seq MOSTA
Methods Numerical Fidelity Statistical Correlation Bio-conservation
[…]
Nearest Interp. 6.24 1.11 0.741 0.695 0.353 0.280
FFN [48] 6.51 1.20 0.706 0.718 0.400 0.143
SIREN [43] 7.21 1.31 0.661 0.678 0.247 0.289
STAGE [21] 6.52 1.11 0.732 0.747 0.365 0.139
SUICA (Ours) 5.66 0.85 0.797 0.792 0.447 0.343
Table 1: Quantitative benchmarking results on MOSTA dataset [2]. Bold figures are best scores and underlined figures are second-best.
Figure 4: Visual comparisons of predicted gene expressions on MOSTA dataset [2].
SUICA outperforms vanilla INR methods [48, 43] and STAGE [21] in predicting the gene expression of unseen spots in the MOSTA dataset [2] as presented in Tab. 1. SUICA not only achieves the best MAE, MSE and cosine similarity, exceeding the second-best method by 5.6%, but also achieves higher correlation, with a significant gap ahead by at least 4.5%. Apart from the reconstruction quality, SUICA surprisingly manages to strengthen the biological signals, by obtaining a bio-conservation score 5.4% higher than the runner-up method and even 3.9% above the ARI calculated with the ground truth.
Fig. 4 showcases SUICA’s ability to closely predict ground-truth gene expression levels, e.g., for RPLP0 gene that highly involves in the ribosomal function and consistently expresses across various regions of mouse embryo [50, 35], SUICA accurately predicts this uniform expression pattern, while other methods overemphasize it in specific regions. SUICA also accurately localizes AFP expression to the liver in the mouse embryo [19]. Additionally, SUICA is the only method to capture the expression of RP58, a transcription factor crucial for mouse brain development [34].
Beyond accurately predicting gene expression in unobserved regions, SUICA is also capable for imputation, enhancing the data to better reflect true underlying biological signatures. For example, despite its low expression level in the ground-truth, SUICA successfully imputes SEPT3, a gene involved in neuronal development, effectively restoring this signal in the brain region. A similar restoration is observed in ELAVL3 of the E10.5 embryo [8]. These results highlight SUICA’s capacity not only to interpolate but also to enrich underlying biological signatures, making it a useful tool for imputing and enhancing spatial gene expression data with high fidelity preserved.
Human and Mouse Brains Visium
Methods Visium-Human Brain Visium-Mouse Brain
[…]
Nearest Interp. 5.36 0.824 0.781 0.365 5.45 0.870 0.736 0.566 0.292
FFN [48] 5.76 0.772 0.786 0.402 5.95 0.832 0.741 0.581 0.000587
SIREN [43] 6.58 0.756 0.747 0.196 5.35 0.878 0.804 0.647 0.359
STAGE [21] 6.19 0.795 0.772 0.223 4.55 0.918 0.825 0.666 0.140
SUICA (Ours) 4.99 0.860 0.846 0.445 3.68 0.932 0.800 0.660 0.393
Table 2: Quantitative benchmarking results on two Visium ST cases [36, 53]. Note that for Visium-Human Brain, there is no annotation of cell type for the evaluation of ARI. Bold figures are best scores and underlined figures are second-best.
Different from Stereo-seq datasets, Visium is a lower-resolution but more affordable ST technology. Tab. 2 shows that SUICA outperforms other methods on the Visium-Human Brain dataset [53]. Due to the lack of independent cell-type annotations, ARI is not applicable in this data. On the Visium-Mouse Brain dataset [36], SUICA demonstrates improved ARI, which indicates that the predictions can more accurately describe the underlying cellular heterogeneity.
Mouse hippocampus Slide-seqV2
Slide-seqV2 allows to sequence RNAs with a spatial resolution of 10 $mm$. The Mouse hippocampus dataset [44] is applied to evaluate the effectiveness of SUICA. As shown in Tab. 3, SUICA achieves a substantially lower MAE and a notably higher cosine similarity (marking a 16.5% improvement) compared to other methods. Fig. 5 illustrates the ground-truth and predicted gene expression from the benchmarking methods. Like other methods, SUICA accurately predicts the expression of hippocampus marker genes, such as HPCA (Hippocalcin) [39] and MALAT1, a gene abundantly expressed in the mouse brain [39]. SUICA also demonstrates its ability to impute the underlying biological signals, as evidenced by GABRB3, which is consistently and strongly expressed in the mouse hippocampus [47]. These findings suggest that SUICA provides reliable predictions that closely reflect ground-truth data, capturing key gene expression patterns in the Slide-seqV2 platforms.
Figure 5: Visual comparisons of the predicted marker gene expression patterns in the Slide-seqV2 Mouse hippocampus dataset [44].
[…]
Nearest Interp. 0.360 0.555 0.361 0.222 0.111
FFN [48] 0.378 0.499 0.442 0.274 0.0523
SIREN [43] 0.383 0.494 0.452 0.248 0.110
STAGE [21] 0.351 0.587 0.483 0.314 0.0361
SUICA (Ours) 0.265 0.752 0.473 0.308 0.111
Table 3: Quantitative results on Slide-seqV2 Mouse hippocampus dataset [44]. Bold figures are best scores and underlined figures are second-best.
4.4 Bio-conservation Analysis of Predicted Cell Type Clusters
Biological variance conservation (Bio-conservation) refers to the degree to which a computational method or model preserves biologically meaningful features, such as cell types, gene expression patterns, or cellular relationships, relative to the ground-truth data [26]. A high bio-conservation score indicates that the predicted data aligns closely with biological reality, capturing key aspects of cell identity, functional states, or tissue structure that are essential for accurate analysis.
In addition to quantitative bio-conservation metrics like ARI used in Section 4.3, Fig. 6 provides a visual assessment of bio-conservation by coloring predicted spots according to dominant cell types in individual clusters. For a fair comparison, we ensured that the number of clusters matched the number of cell-type labels in the dataset.
In the mouse embryo E16.5, apart from the highest ARI among benchmarking methods, SUICA also offers the closest spatial visualization of cell types to the ground truth. Compared to other methods, SUICA captures fine-grained structures, such as the choroid plexus in the brain, the gastrointestinal (GI) tract, and the muscle cells surrounding the heart. FFN’s [48] predictions appear chaotic, with most cell-type predictions misaligned with the ground-truth labels. While SIREN [43] correctly captures general cell types patterns, it overlooks many details detected by SUICA, such as the connective tissue around the cartilage primordium in the mouse’s foot region and the cavity surrounding the heart. STAGE [21] identifies several major cell types in the mouse embryo—such as liver and brain—but misses smaller, yet important, cell populations.
In the Mouse hippocampus dataset, SUICA also demonstrates its ability to capture both global and fine-grained features, accurately identifying endothelial cells, Ependymal cells, and inter-neurons. In contrast, FFN and STAGE fail to detect smaller cell populations, while SIREN misclassifies inter-neurons. Similarly, on the Visium-Mouse Brain dataset, SUICA uniquely identifies specific cell populations, such as Lateral Ventricle cells and Cortex_5, which are absent from the predictions of other methods.
Figure 6: Spatially visualized comparison on bio-conservations of predicted spots on MOSTA mouse embryo E16.5, Slide-seqV2 mouse hippocampus, and Visium-Mouse brain.
4.5 Ablation Study
We showcase two examples to illustrate how the proposed modules in SUICA contribute to the final model performance, namely the E16.5 embryo of Stereo-seq MOSTA (spatially dense, 121,756 cells, 26,159 genes, with FFN [48] as backbone) and the Human Brain Visium (spatially sparse, 4,910 cells, 36,592 genes, with SIREN [43] as backbone).
Settings Embryo E16.5 Human Brain
[…]
Vanilla INR 2.35 0.668 0.653 9.33 0.756 0.747
+AE 1.60 0.789 0.751 11.27 0.695 0.691
+Dice 1.48 0.806 0.747 7.05 0.826 0.800
     +Graph 1.47 0.807 0.761 5.67 0.860 0.846
Table 4: Ablation study on the key modules in SUICA.
As can be indicated from Tab. 4, the spatial density and sequencing depth may influence SUICA’s effectiveness. For the Embryo E16.5 dataset, the primary performance improvement comes from the AE and the quasi-classification loss. In contrast, in the Human Brain dataset, adding AE causes a performance decrease, while the GCN has a more significant impact on overall performance. Intuitively, we attribute this difference to the varying spatial sparsity of different ST techniques and GAE alleviates such issue by incorporating spatial context.
5 Conclusion
In this paper, we seek to model the super-high dimensional and sparse nature of ST data, enhancing both spatial resolution and gene expression with the smooth prior inherent in INRs. To this end, we present SUICA, a powerful INR variant tailored to model ST in a continuous and compact manner. Using a Graph Autoencoder, SUICA maps zero-inflated raw data into a lower-dimensional embedding space, preserving high-frequency details and making the complex embedding mapping more feasible for INRs. The INR fitted embeddings are then decoded to the raw expression with the decoding head preventing the prediction error from accumulating. To encourage the sparsity in the final predictions, we leverage a quasi-classification loss term as a regularizer, preserving both the sparsity and numerical fidelity of non-zero values. Extensive experiments on Stereo-seq, 10x Genomics Visium, and Slide-seqV2 datasets demonstrate SUICA’s effectiveness, yielding improvements in both in-silico metrics and biologically meaningful analyses, with enhanced spatial resolution and biological signatures. It is firmly believed by us that SUICA will be an inspiring attempt from both INRs and ST perspectives.
Supplementary Material
Due to the page limit of the main text, we have included some background information, some discussions regarding the model choices in SUICA, and some other experimental results we find meaningful for the readers.
6 Background
Spatial Transcriptomics (ST) is a technique that representing the gene expression at individual spots within a slice of sample, providing a detailed picture of cellular activity and composition in the micro-environment. ST sequencing is typically performed on fresh frozen tissue samples and is broadly categorized into two main types: sequencing-based ST and imaging-based ST. Despite differences in methodology, both technologies deliver complementary measurements of in situ gene expression. The sequencing-based ST technologies, including Stereo-seq, 10x Visium, Slide-seqV2, usually rely on spatially indexed surfaces to encode spatial information (Supplementary Fig. 7). These methods employ different spatial indexing strategies: 10x Visium uses a microarrayer spotting robot, these spots are 50–200 $\mu m$ in size, to deliver a unique barcode to a fixed, known location on the surface of a slide [51]; Slide-seq applies 10 $\mu m$ diameter nanobeads for spatial barcoding [51]; Stereo-seq uses DNA nanoball-patterned arrays for in situ RNA capture to create high-resolution spatial transcriptomics sequencing [2]. Following barcoding, mRNA diffuses to spatially indexed primers, undergoes reverse transcription, cDNA amplification, and short-read sequencing. This process generates transcript reads paired with barcode sequences, localizing each detected transcript to specific pixels.
For the imaging based-ST methods, RNA molecules are specifically tagged with fluorescent probes by complementary hybridization. These probes are then imaged using fluorescence microscopy. MERFISH, for instance, maps RNA molecules with binary barcodes encoded with error-correcting codes to ensure accuracy. Sequential rounds of fluorescence imaging detect the presence or absence of fluorescence, decoding the RNA identity and precisely mapping molecules to their spatial locations.
Figure 7: Workflow of sequencing-based spatial transcriptomics sequencing. (a) Spatially indexing, with microarrayer for 10x Visium, nanobead technology for Slide-seqV2 and nanoball-based strategy for Stereo-seq; (b) Placing the slice of tissue on the indexed surface and capturing mRNA from each spot; (c) Collecting spots and conducting sequencing; (d) The generated data matrix with gene expression and location information for each spot.
7 Model Details & Insights
FFN vs. SIREN
As mentioned in the main body, it is observed that when applied to ST with low spatial density, SIREN [43] outperforms FFN [48] as the backbone of SUICA, while for high-resolution ST slices, FFN prevails. To better showcase such phenomenon, we provide two representative cases of the mouse embryos E9.5 and E16.5 in Supplementary Tab. 5. We globally set the angular velocity $\omega$ of the periodical activation function in SIREN [43] as 30 which is adopted by default and the mapping size of the Random Fourier Features as 256. For more information for these two magic numbers, please refer to the respective papers. Note that we do not intend to showcase any superiority among INR variants. Since we adopt a modular design that is agnostic to specific INR implementations, we are able to take advantage of different implementations to meet varying requirements.
Backbones Embryo E9.5 Embryo E16.5
[…]
FFN [48] 0.55 0.752 0.799 1.47 0.807 0.761
SIREN [43] 0.52 0.766 0.814 1.62 0.786 0.745
Table 5: Quantitative results of SUICA empowered by different INR backbones.
Explicit Alternatives
Recently, researchers are employing explicit or hybrid alternatives [32, 3] to replace the fully implicit representations [48, 43, 23] for a faster inference speed. For a more comprehensive benchmarking, we here compare SUICA with the trainable hash-based encoding [32] (denoted as NGP) on Stereo-seq MOSTA dataset [2]. The quantitative results are shown in Supplementary Tab. 6. For the implementation, we re-compile the library of tiny-cuda-nn to support float32, to align with other methods.
[…]
NGP [32] 6.39 1.14 0.729 0.742 0.415
SUICA 5.66 0.85 0.797 0.792 0.447
Table 6: Quantitative comparison on Stereo-seq MOSTA dataset [2] between SUICA and NGP [32].
CNN
It is easy to notice that the network architecture in SUICA is almost fully MLP-based except for the GCN. Such architecture does not explicitly involve any local inductive bias that CNNs usually do. The underlying reason for doing so is that along the sequencing depth of ST, it is assumed that the channels (i.e., specific gene segments) are permutation-invariant even though there might be some statistical correlations. In this way, it is not supposed to make use of the context information along the sequencing depth but for the spatial adjacency, we model such context by the GCN.
PCA
Note that by aid of GAE, we reduce the dimension of raw gene expressions from thousands to 32, which is recovered later by the decoder. The job of the GAE can be replaced by PCA (Principal Component Analysis), by a pre-fitted linear projection. However, it is obviously challenging for the linear transformation of PCA and inverse PCA to model the complex distribution of ST, leading to a great loss in the reconstructed signal. Empirically, it is observed that when encoding and decoding with PCA and inverse PCA, the IoU between the predicted zero maps and the ground truth decreases to 0, indicating a complete failure of recovering the sparsity, while SUICA maintains an IoU of around 0.9.
VAE
Variational Autoencoders (VAEs) are a common alternative to AEs, designed to generate samples from a probabilistic latent space, typically modeled as Gaussian distributions. While VAEs offer generative capabilities through variational inference and learn smoother, more continuous latent spaces compared to AEs, they often produce blurry outputs [60] and struggle to match AEs in reconstruction accuracy [7]. These issues might lead to ambiguity in gene expression that is not wanted. These limitations stem from the inherent trade-off between the quality of latent representation learning and reconstruction fidelity in VAEs. Although variants like beta-VAE [13] allow for adjustable weighting between these objectives, the issue persists. In our setting, where numerical fidelity and reconstruction accuracy take precedence over generative and latent representation quality, we select AE over VAE for SUICA’s design, leveraging its ability to accurately reconstruct high-dimensional raw gene expression data.
Reconstruction Loss
In SUICA, we train the INR and the decoder separately to minimize the accumulated prediction error, on which we would like to further offer some insights. Obviously, a pre-trained decoder is necessary; otherwise the decoder would become a trivial extension of the INR. It is also easy to understand that such decoder requires finetuning to bridge the gap between the pseudo ground-truth embeddings $\mathbf{z}_{\text{gt}}$ and the predicted embeddings $\hat{\mathbf{z}}$. The problem that prevents this two-stage training strategy from merging into one end-to-end stage is that when using $\mathcal{L}_{\text{recons}}$ through the decoder to supervise the INR, the loss will become unstable. We attribute this observation to the optimizing objectives of $\mathcal{L}_{\text{embd}}$ and $\mathcal{L}_{\text{recons}}$ are somehow contradictory. Though this problem can be alleviated by carefully adjusting the training-relevant hyper-parameters, we opt to adopt the two-stage training strategy for simplicity and stable performance.
8 STAGE
In the main text, we involve STAGE [21] for both quantitative and qualitative comparisons as a baseline for histology-free super-resolution of gene expressions. In addition to solely showing the superiority of SUICA, we would like to shed some light on this matter.
Regarding the benchmarking task of super-resolution, we identify two critical sub-tasks, namely the reconstruction of gene expression and the spatial mapping. STAGE models the super-resolution as a generative task, where an AE is adopted to self-regress the raw representations while enforcing the latent as the corresponding 2D coordinates. In this way, the two sub-tasks are coupled as one. Considering the challenges posed, the closely coupled paradigm may not lead to satisfactory results. On the contrary in SUICA, the two sub-tasks are well decoupled, where the spatial mapping is performed by the INR while the description as well as the reconstruction is taken care of the GAE. The progressive training paradigm also guarantees that each module is doing the assigned job, keeping the coupling at minimum.
According to the experimental results, SUICA not only achieves better quantitative scores for benchmarking, but also exhibits much richer bio-conservation for downstream applications.
9 MERFISH
MERFISH is an imaging-based ST technique that enables the highly multiplexed imaging of RNA molecules in cells while maintaining their spatial context. Compared to the Slide-seqV2, 10x Visium, and Stereo-seq technologies used in our manuscript, MERFISH data can quantify significantly fewer genes for individual cells. In Supplementary Tab. 7, we evaluate the performance of SUICA using a human heart MERFISH dataset [10] with 228,635 cells and 238 genes. SUICA achieves a significant lower mean absolute error (MAE) and mean square error (MSE), and the cosine similarity is at least 2.28% higher than other methods, showing its high numerical fidelity. Following the bio-conservation analysis scheme, we annotate the cell types for predicted spots of MERFISH data using dominant cell types in individual clusters. We visualize the spatial bio-conservation by coloring the cell types in Supplementary Fig. 8. The results show that the SUICA predictions mimic the ground-truth evidenced the capability of SUICA to predict unseen spots in the MERFISH datasets.
Figure 8: Spatially visualized predicted spots on MERFISH human heart [10].
[…]
Nearest Interp. 6.43 6.53 0.814 0.641 0.476
FFN [48] 6.06 5.86 0.840 0.717 0.554
SIREN [43] 5.54 4.98 0.864 0.759 0.606
STAGE [21] 5.48 5.10 0.870 0.709 0.558
SUICA 4.65 3.92 0.892 0.718 0.548
Table 7: Quantitative results on MERFISH human heart [10].
10 Data Pre-processing
To enable the evaluation of Pearson Correlation Coefficient and Spearman’s Rank Correlation Coefficient, before the experiments we remove empty rows and lines to make sure each spot is at least expressed by a limited number of genes.
Note that the outputs of conventional INR-based tasks are strictly bounded, where the linear layer is usually clamped or compressed by sigmoid. For ST data, we do not have such assumption and only employ ReLU as the final activation. Therefore, as a common practice in ST analysis [28, 17], we remove the genes whose expressions are overly high and normalize each cell by total counts over all genes of the cell so that every cell has the same total count after normalization while keeping the original sparsity.
11 Limitations
One of the significant limitations in SUICA roots in the case-by-case training paradigm of INRs. Similar to NeRF [29], for each case, i.e., each ST slice, we need to training a new model from scratch. Considering the overwhelming heterogeneity of different ST data, we temporarily compromise in this issue but incorporate domain knowledge to make SUICA generalizable is an interesting future work.
Another potential limitation is that SUICA prefers high-quality ST data, in terms of both spatial density and sequencing depth, as can be indicated from the experimental results we provide. SUICA exhibits the most performance gain in Stereo-seq compared to other platforms and when the gene expressions no longer meet the super-high dimensional sparse assumption, SUICA will degenerate to a vanilla INR, as is consistent with the results of MERFISH in Supplementary Tab. 7.